OpenStudy (anonymous):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 3/4?

OpenStudy (anonymous):

\[y=mx+b\]

OpenStudy (anonymous):

@tylermcmullen23 do you think you can show me how to set it up with the numbers i got because i dont understand how to?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

the slope given goes in place of the "m" variable. so now y=3/4x+b

OpenStudy (anonymous):

in order to find b (y intercept) you have to use point slope form and solve for y. \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

replace y_1 with 4 and m with 3/4 and x_1 with 5 so that now, y+4=3/4x-5

OpenStudy (anonymous):

subtract the 4 from both sides to get y by itself: y=3/4x-9